# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 5.1 Pipeline performance
Figure 13: Bar chart comparing the run-time of a single evaluation of KiDS-SBI (above) versus a single evaluation of a simulation based on Joachimi et al. (2021) (below), both on a single core ($N_{\mathrm{shells}}=19$, $N_{\mathrm{tomo}}=5$ and $N_{\mathrm{side}}=1024$). Both suites of simulations use CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012) to compute the three-dimensional matter power spectrum. For the reference simulations, we use the non-Limber projection built into CAMB with limber_phi_lmin $=1200$ rather than Levin with $\ell_{\mathrm{max,nL}}=1200$. We run FLASK (Xavier et al., 2016) rather than GLASS (Tessore et al., 2023) to compute the underlying matter and convergence fields of each of the 19 shells. Subsequently, we sample galaxies using SALMO in both cases, and then calculate the spatial two-point correlation functions, $\xi_{\pm}(\theta)$, rather than calculating the angular power spectra, $\tilde{C}(\ell)$. To calculate $\xi_{\pm}(\theta)$ in the reference simulations, we use TreeCorr (Jarvis et al., 2004).
@@ 5.2 Parameters and priors
To avoid overinformative priors, most are flat top-hat functions. For the same reason, the top-hat priors in $S_{8}$, $h_{0}$, $\omega_{\mathrm{b}}$, $A_{\mathrm{IA}}$ and $A_{\mathrm{bary}}$ are chosen to be wide. The prior on $n_{\mathrm{s}}$ spans a smaller range around the theoretical value of unity for scale-invariant primordial fluctuations. This avoids artefacts within the prior volume as this parameter is not well constrained by weak gravitational lensing. The prior on $\omega_{\mathrm{c}}$ is defined to be consistent with a range in $\Omega_{\mathrm{m}}\in[0.188,0.408]$, where the limits are given by the ±$\pm 5\sigma$ intervals of the marginal constraints from independent measurements of luminosity distance to Type Ia Supernovae in Scolnic et al. (2018). See Joachimi et al. (2021) for more details on the motivations for the chosen priors. The priors on the nuisance parameters capturing any shifts in the mean of each of the five tomographic bins, $\bm{\delta}_{z}$, are given by a multivariate Gaussian, since the $\delta_{z}$ of a given tomographic bin is not independent of the shifts in the other bins. This is quantified by the covariance, $\bm{C}_{z}$, as estimated in Hildebrandt et al. (2021).
@@
Table 2: Table showing the parameters which are varied within the simulation-based inference pipeline. The prior ranges are selected to be exactly in line with previous KiDS-1000 analyses (Asgari et al., 2021; Heymans et al., 2021; Loureiro et al., 2022; van den Busch et al., 2022; Tröster et al., 2022). The upper five rows show the cosmological parameters of interest, while the three rows below show the nuisance parameters which quantify systematic biases. The last three rows show shear bias parameters for each tomographic bin, $p$, as given in Fig. 5 which are sampled implicitly within each simulation by drawing from the prior distribution, so any posterior is pre-marginalised over these parameters. For flat priors, the lower and upper limits of the normalised rectangular function defines the prior. For the Gaussian prior on $\bm{\delta}_{z}$, we use a five-dimensional multivariate Gaussian with its mean at the zero vector and the covariance, $\bm{C}_{z}$, defined by the one estimated in (Hildebrandt et al., 2021). Note that for simplicity the $\bm{\delta}_{z}$ are implicitly marginalised throughout this analysis. The fiducial values given here constitute the fiducial parameter choice for the score compression described in Sect. 5.3.
@@ 5.3 Score compression
If the likelihood is known a priori, one can compress a given data vector to a summary of the same dimensionality as the degrees of freedom in the assumed model, such that Fisher information is conserved (Zablocki & Dodelson, 2016; Alsing & Wandelt, 2018; Alsing et al., 2018, 2019). This allows us to compress the data vector, $\bm{d}$, down to a vector, $\bm{t}$, with the dimensionality of $|\bm{\Theta}|$ as follows
@@
Additionally, we find that the 12 parameters shown in Table 2 are not necessary to characterise the degrees of freedom of the 120-dimensional cosmic shear pseudo-Cls. As all $\delta_{z}$ are broadly consistent with zero, we may compress all data vectors to a seven-dimensional summary which is still capable of capturing all the complexity in the data. In any case, the $\bm{\delta}_{z}$ parameters are still explicitly varied within the simulations to propagate any uncertainties on the mean of the tomographic bins. For simplicity, henceforth all posteriors are implicitly marginalised over the five $\delta_{z}$ parameters.
@@ 5.4 Density estimation likelihood-free inference (DELFI)
As the first step in the exploration of the prior space, the ensemble of these neural density estimators (NDEs) initially learns the sampling distribution from a set 2000 points on a Latin hypercube. Based on this, we then make use of the active learning feature within DELFI to sample additional points in parameter space which efficiently contribute to learning the effective likelihood within prior space. To ensure that the sampling distribution density, $P(\bm{t}|\bm{\Theta},\bm{w})$ is well learnt, recursive sampling continues until the validation loss stops decreasing for 20 training epochs, where the loss function is defined as,
@@
We find that for the anisotropic systematics model, this occurs already after 5,000 realisations. We note that the NDEs require fewer realistic KiDS-SBI forward simulations to minimise the loss function as is the case in L23 where the forward simulations were idealised random samples from a Gaussian distribution. In any case, to ensure that the prior space is densely sampled, we choose to train the NDEs on 18,000 realisations. This is facilitated by the computational efficiency of the forward simulations making additional realisations relatively inexpensive.
@@
As advised in L23, to minimise the loss of information through from the compression step described in Sect. 5.3, once the simulation-based inference has been conducted the first time on the KiDS-1000 data for a given forward-model, we repeat the inference again a second time while taking the fiducial model parameters assumed in the score compression to be given by the maximum a posteriori (MAP) from the initial inference (see Tab. 3). Note that in the second inference, we do not change the forward model in any way, and keep the exact same realisations of the model as in the first analysis. Only the fiducial data vector used to characterise the likelihood in Eq. (43) are updated to re-compress all simulated data vectors. Considering that the initial fiducial parameters based on previous KiDS-1000 analyses are mostly consistent with the MAP which we infer in the initial analysis, while score compression is also robust against the biased fiducial parameters (Lin et al., 2023), we find that this step only marginally alters the inferred posterior.
@@ 6 Goodness-of-fit in SBI
For our setup, we evaluate a mean model data vector and a covariance from a set of 1,600 simulations run at the fiducial cosmology. It should be noted that these simulations are the same as the ones required for the score compression step described in Sect. 5.3, so this comes without any additional computational costs. Using this newly obtained model data vector and covariance, and the $\chi^{2}$ discrepancy measure defined above, we calculate the discrepancy measure for the 100 separate forward simulations from the anisotropic systematics model, and the KiDS-1000 gold sample data, which are depicted in the left panel of Fig. 14. As we obtain a probability to exceed (PTE) of 0.42, we find the fit for the data to be good.
Note that, in this test, the degrees of freedom are not recovered directly due to the effects of normalisation and optimal score compression in doing dimensionality reduction.
@@ 7.1 Coverage
Refer to captionRefer to caption
Figure 16: Left panel: Comparison of marginalised posterior contours from noise-free data assuming the anisotropic systematics model (which includes variable depth and PSF variations) in blue and green, and the standard isotropic systematics model (which is in line with the analysis assumptions in Asgari et al. 2021, i.e. it does not model variable depth or PSF variations) in orange and pink. The orange and the blue posteriors are derived from the same data vector: a noise-free realisation from the anisotropic systematics model at the fiducial cosmology shown in Table 2; while the pink and the green posteriors are derived from a noise-free realisation from the standard isotropic systematics model at the same cosmology. Right panel: Comparison of marginalised posterior contours from noisy data simulated with anisotropic systematics model with the same seed as in the noise-free case, and subsequently analysed with the SBI based on the anisotropic systematics model in blue, and standard isotropic systematics model in orange. In both panels, the black solid line indicates the true cosmology assumed for the mock data vector (see Table 4).
To determine whether the learnt posterior distribution from the anisotropic systematics model is representative of the true underlying posterior and is unbiased, we conduct the Tests of Accuracy with Random Points (TARP) shown in Fig. 15. TARP (Lemos et al., 2023a) measures the expected coverage probability of random posterior samples within a given credibility level of the learnt posterior empirically. Whilst the original authors have their implementation of the coverage test, we made use of our implementation (ECP-GF17) to interface with the outputs from our simulations and DELFI in a more straightforward manner.
@@
We then used our trained DELFI model to perform inference on the new suite of 500 simulations, each acting as a mock data vector and thus obtaining 1000 posterior points for all 500 simulations. From the posterior samples, the coverage probability was calculated empirically for a given credibility level (see Lemos et al. 2023a for a detailed discussion of the algorithm). Due to the inherently stochastic nature of the coverage probability test, the test was bootstrapped 100 times for each credibility level. Fig. 15 depicts both the mean and two standard deviations of the spread of the bootstrapped TARP coverage test.
@@ 7.2 Impact of variable depth and shear bias
To assess the impact of the additional observational systematics considered in the anisotropic systematics model, we re-train a set of NDE ensembles on another 18,000 simulations of a different model which does not consider variable depth or additional shear biases while following the same procedure as outlined in Sect. 5. We refer to this model as the standard isotropic systematics model as it is more in line with previous ”standard” KiDS-1000 analyses such as Asgari et al. (2021), van den Busch et al. (2022), Li et al. (2023a), or Dark Energy Survey and Kilo-Degree Survey Collaboration et al. (2023). This model considers the cosmological signal, intrinsic alignments, and baryonic feedback in the same way as the anisotropic systematics model, but it has two main differences: it does not consider variable depth, and it does not consider the additive or PSF shear biases. Firstly, when sampling galaxies as described in Eq. (20) the galaxy density, $n_{\mathrm{gal}}$, and the photometric redshift distribution, $P(p|i)$, are assumed to be isotropic for a given tomographic bin. Secondly, when biasing the observed galaxy shapes according to the shear biases measured in KiDS-1000, the standard isotropic systematics model only considers the multiplicative shear bias, i.e. only $M^{(p)}\neq 0$ while $\alpha^{(p)}=0,\,c^{(p)}=0\,\forall p$.
This simplification of the forward model impacts the measured cosmic shear signal in a few ways. Firstly, from just comparing the measured two-point statistics, as can be seen in more detail in Appendix C, we find that variable depth can contribute up to $1\%$ to the measured two-point shear signal in KiDS-1000, which is consistent with the predictions from Heydenreich et al. (2020). At the same time, variable depth can significantly alter the uncertainty of the two-point statistic (by $\lesssim 20\%$; Joachimi et al. 2021). As the intrinsic galaxy shape dispersion and the galaxy density vary anisotropically, the shape noise of the sample is no longer determined by the mean value per tomographic bin. Instead, the shape noise becomes an anisotropic distribution too which in the case of KiDS-1000 is skewed below the mean (see Appendix C). In addition, as can be seen by eye from Fig. 7 and 8, variable depth can add angular correlations to the data at the scales of pixels, pointing overlaps, whole pointings or even over the whole footprint, while also changing the signal along the line-of-sight anisotropically as shown in Fig. 10. These additional correlations can significantly alter the covariance of the two-point statistic or may even lead to non-Gaussian noise.
@@
Figure 17: Comparison of likelihood models with a noise-free mock data vector from the anisotropic systematics forward model. Shown are marginalised posterior contours: in blue, assuming the anisotropic systematics model within the standard setup of KiDS-SBI as described in Sect. 5.4; in orange, assuming the same anisotropic systematics forward simulations to train a Mixture Density Network (MDN; Bishop 1994) which is made up of a single multivariate Gaussian; and in pink, applying a standard analysis which assumes a Gaussian likelihood characterised by a numerical sample covariance derived from anisotropic systematics forward simulations and a model given by the analytic pseudo-Cl model described in Appendix D. All posteriors are evaluated assuming the same noise realisation and the same model parameters (see Table 4).
The other distinguishing systematic between the anisotropic systematics model and the standard isotropic systematics model is the shear bias. When including additive and PSF shear biases in the anisotropic systematics model, that can also add a per-cent level (up to ${\sim}0.3\sigma$, but less than $0.1\sigma$ at most angular scales) non-cosmological signal to the pseudo-Cls, mostly driven by the PSF variation shear bias (see Appendix C for details). This is consistent with the negligible impact on the measurement found in previous studies (Giblin et al., 2021). At the same time, the inclusion of the $\alpha_{1}^{(p)}$, $\alpha_{2}^{(p)}$, $c_{1}^{(p)}$ and $c_{2}^{(p)}$ parameters for each tomographic bin and their associated uncertainties adds additional variance to the anisotropic systematics model when we pre-marginalise over these 20 parameters. Therefore, the additive and PSF variation shear biases are yet another source of additional non-cosmological signal and variance.
@@
To assess whether the inclusion or exclusion of variable depth and shear biases can bias the SBI of cosmological parameters, we repeat the inference of a mock data vector, but this time we generate a random realisation from the standard isotropic systematics model and analyse it using the likelihood learnt from the forward simulation which assumes the anisotropic systematics model, and vice versa. This allows us to determine whether the inclusion of the effects in the forward model would bias our inference if they were not actually present in the data that is analysed. We also show these posteriors in Fig. 16 where we see that when the data model and the likelihood model are mismatched, the constraints are consistent within a given model. This confirms that, given the learned likelihood, the bias in the two-point shear signal from both variable depth and PSF shear bias is negligible in KiDS-1000. These findings are consistent with previous work which found this to be the case for KiDS-1000 for variable depth (Heydenreich et al., 2020; Joachimi et al., 2021) and for the PSF shear bias (Giblin et al., 2021).

### v2

@@ 5.1 Pipeline performance
Figure 12: Bar chart comparing the run-time of a single evaluation of KiDS-SBI (above) versus a single evaluation of a simulation based on Joachimi et al. (2021) (below), both on a single core ($N_{\mathrm{shells}}=19$, $N_{\mathrm{tomo}}=5$ and $N_{\mathrm{side}}=1024$). Both suites of simulations use CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012) to compute the 3D matter power spectrum. For the reference simulations, we use the non-Limber projection built into CAMB with limber_phi_lmin $=1200$ rather than Levin with $\ell_{\mathrm{max,nL}}=1200$. We run FLASK (Xavier et al., 2016) rather than GLASS (Tessore et al., 2023) to compute the underlying matter and convergence fields of each of the 19 shells. Subsequently, we sample galaxies using SALMO in both cases, and then calculate the spatial two-point correlation functions, $\xi_{\pm}(\theta)$, rather than calculating the angular power spectra, $\tilde{C}(\ell)$. To calculate $\xi_{\pm}(\theta)$ in the reference simulations, we use TreeCorr (Jarvis et al., 2004).
@@ 5.2 Parameters and priors
To avoid overinformative priors, most are flat top-hat functions. For the same reason, the top-hat priors in $S_{8}$, $h_{0}$, $\omega_{\mathrm{b}}$, $A_{\mathrm{IA}}$, and $A_{\mathrm{bary}}$ are chosen to be wide. The prior on $n_{\mathrm{s}}$ spans a smaller range around the theoretical value of unity for scale-invariant primordial fluctuations. This avoids artefacts within the prior volume as this parameter is not well constrained by weak gravitational lensing. The prior on $\omega_{\mathrm{c}}$ is defined to be consistent with a range in $\Omega_{\mathrm{m}}\in[0.188,0.408]$, where the limits are given by the ±$\pm 5\sigma$ intervals of the marginal constraints from independent measurements of luminosity distance to Type Ia Supernovae in Scolnic et al. (2018). See Joachimi et al. (2021) for more details on the motivations for the chosen priors. The priors on the nuisance parameters capturing any shifts in the mean of each of the five tomographic bins, $\bm{\delta}_{z}$, are given by a multivariate Gaussian, since the $\delta_{z}$ of a given tomographic bin is not independent of the shifts in the other bins. This is quantified by the covariance, $\bm{C}_{z}$, as estimated in Hildebrandt et al. (2021).
Table 2: Parameters that are varied within the simulation-based inference pipeline and their priors.
@@
 [^18] The prior ranges are selected to be exactly in line with previous KiDS-1000 analyses (Asgari et al., 2021; Heymans et al., 2021; Loureiro et al., 2022; van den Busch et al., 2022; Tröster et al., 2022). The upper five rows show the cosmological parameters of interest, while the three rows below show the nuisance parameters which quantify systematic biases. The last three rows show shear bias parameters for each tomographic bin, $p$, as given in Fig. 5 which are sampled implicitly within each simulation by drawing from the prior distribution, so any posterior is pre-marginalised over these parameters. For flat priors, the lower and upper limits of the normalised rectangular function defines the prior. For the Gaussian prior on $\bm{\delta}_{z}$, we use a 5D multivariate Gaussian with its mean at the zero vector and the covariance, $\bm{C}_{z}$, defined by the one estimated in (Hildebrandt et al., 2021). We note that for simplicity, the $\bm{\delta}_{z}$ are implicitly marginalised throughout this analysis. The fiducial values given here constitute the fiducial parameter choice for the score compression described in Sect. 5.3.
@@ 5.3 Score compression
If the likelihood is known a priori, we can compress a given data vector to a summary of the same dimensionality as the degrees of freedom in the assumed model, such that Fisher information is conserved (Zablocki & Dodelson, 2016; Alsing & Wandelt, 2018; Alsing et al., 2018, 2019). This allows us to compress the data vector, $\bm{d}$, down to a vector, $\bm{t}$, with the dimensionality of $|\bm{\Theta}|$ as follows
@@
Additionally, we find that the 12 parameters shown in Table 18 are not necessary to characterise the degrees of freedom of the 120-dimensional cosmic shear pseudo-Cls. As all $\delta_{z}$ are broadly consistent with zero, we may compress all data vectors to a 7D summary which is still capable of capturing all the complexity in the data. In any case, the $\bm{\delta}_{z}$ parameters are still explicitly varied within the simulations to propagate any uncertainties on the mean of the tomographic bins. For simplicity, henceforth all posteriors are implicitly marginalised over the five $\delta_{z}$ parameters.
@@ 5.4 Density estimation likelihood-free inference (DELFI)
As the first step in the exploration of the prior space, the ensemble of these neural density estimators (NDEs) initially learns the sampling distribution from a set 2000 points on a Latin hypercube. Based on this, we then make use of the active learning feature within DELFI to sample additional points in parameter space which efficiently contribute to learning the effective likelihood within prior space. To ensure that the sampling distribution density, $P(\bm{t}|\bm{\Theta},\bm{w})$ is well learnt, recursive sampling continues until the validation loss stops decreasing for 20 training epochs, where the loss function is defined as
@@
We find that for the anisotropic systematics model, this occurs already after 5,000 realisations. We note that the NDEs require fewer realistic KiDS-SBI forward simulations to minimise the loss function as is the case in L23 where the forward simulations were idealised random samples from a Gaussian distribution. In any case, to ensure that the prior space is densely sampled, we chose to train the NDEs on 18,000 realisations. This is facilitated by the computational efficiency of the forward simulations making additional realisations relatively inexpensive.
@@
As advised in L23, to minimise the loss of information through from the compression step described in Sect. 5.3, once the simulation-based inference has been conducted the first time on the KiDS-1000 data for a given forward model, we repeat the inference again a second time while taking the fiducial model parameters assumed in the score compression to be given by the maximum a posteriori (MAP) from the initial inference (see Tab. 20). We note that in the second inference, we do not change the forward model in any way, and keep the exact same realisations of the model as in the first analysis. Only the fiducial data vector used to characterise the likelihood in Eq. (43) are updated to re-compress all simulated data vectors. Considering that the initial fiducial parameters based on previous KiDS-1000 analyses are mostly consistent with the MAP which we infer in the initial analysis, while score compression is also robust against the biased fiducial parameters (Lin et al., 2023), we find that this step only marginally alters the inferred posterior.
@@ 6 Goodness-of-fit in SBI
For our setup, we evaluate a mean model data vector and a covariance from a set of 1,600 simulations run at the fiducial cosmology. It should be noted that these simulations are the same as the ones required for the score compression step described in Sect. 5.3, so this comes without any additional computational costs. Using this newly obtained model data vector and covariance, and the $\chi^{2}$ discrepancy measure defined above, we calculate the discrepancy measure for the 100 separate forward simulations from the anisotropic systematics model, and the KiDS-1000 gold sample data, which are depicted in the left panel of Fig. 13. As we obtain a probability to exceed (PTE) of 0.42, we find a good fit to the data. We note that in this test, the degrees of freedom are not recovered directly due to the effects of normalisation and optimal score compression in doing dimensionality reduction.
@@ 7.1 Coverage
Figure 15: Contour plots comparing the effects of model misspecification for different model cases and noise levels in the data. Left panel: Comparison of marginalised posterior contours from noise-free data assuming the anisotropic systematics model (which includes variable depth and PSF variations) in blue and green, and the standard isotropic systematics model (which is in line with the analysis assumptions in Asgari et al. 2021, i.e. it does not model variable depth or PSF variations) in orange and pink. The orange and the blue posteriors are derived from the same data vector: a noise-free realisation from the anisotropic systematics model at the fiducial cosmology shown in Table 18; while the pink and the green posteriors are derived from a noise-free realisation from the standard isotropic systematics model at the same cosmology. Right panel: Comparison of marginalised posterior contours from noisy data simulated with anisotropic systematics model with the same seed as in the noise-free case, and subsequently analysed with the SBI based on the anisotropic systematics model in blue, and standard isotropic systematics model in orange. In both panels, the black solid line indicates the true cosmology assumed for the mock data vector (see Table 21).
To determine whether the learnt posterior distribution from the anisotropic systematics model is representative of the true underlying posterior and is unbiased, we conduct the Tests of Accuracy with Random Points (TARP) shown in Fig. 14. TARP (Lemos et al., 2023a) measures the expected coverage probability of random posterior samples within a given credibility level of the learnt posterior empirically. Whilst the original authors have their implementation of the coverage test, we made use of our implementation (ECP-GF19) to interface with the outputs from our simulations and DELFI in a more straightforward manner.
@@
We then used our trained DELFI model to perform inference on the new suite of 500 simulations, each acting as a mock data vector and thus obtaining 1000 posterior points for all 500 simulations. From the posterior samples, the coverage probability was calculated empirically for a given credibility level (see Lemos et al. 2023a for a detailed discussion of the algorithm). Due to the inherently stochastic nature of the coverage probability test, the test was bootstrapped 100 times for each credibility level. Figure 14 depicts both the mean and two standard deviations of the spread of the bootstrapped TARP coverage test.
@@ 7.2 Impact of variable depth and shear bias
To assess the impact of the additional observational systematics considered in the anisotropic systematics model, we re-train a set of NDE ensembles on another 18,000 simulations of a different model which does not consider variable depth or additional shear biases while following the same procedure as outlined in Sect. 5. We refer to this model as the standard isotropic systematics model as it is more in line with previous ’standard’ KiDS-1000 analyses such as Asgari et al. (2021), van den Busch et al. (2022), Li et al. (2023a), or Dark Energy Survey and Kilo-Degree Survey Collaboration et al. (2023). This model considers the cosmological signal, intrinsic alignments, and baryonic feedback in the same way as the anisotropic systematics model, but it has two main differences: it does not consider variable depth, and it does not consider the additive or PSF shear biases. Firstly, when sampling galaxies as described in Eq. (20) the galaxy density, $n_{\mathrm{gal}}$, and the photometric redshift distribution, $P(p|i)$, are assumed to be isotropic for a given tomographic bin. Secondly, when biasing the observed galaxy shapes according to the shear biases measured in KiDS-1000, the standard isotropic systematics model only considers the multiplicative shear bias; that is, only $M^{(p)}\neq 0$ while $\alpha^{(p)}=0,\,c^{(p)}=0\,\forall p$.
This simplification of the forward model impacts the measured cosmic shear signal in a few ways. Firstly, from just comparing the measured two-point statistics, as can be seen in more detail in Appendix C, we find that variable depth can contribute up to $1\%$ to the measured two-point shear signal in KiDS-1000, which is consistent with the predictions from Heydenreich et al. (2020). At the same time, variable depth can significantly alter the uncertainty of the two-point statistic (by $\lesssim 20\%$; Joachimi et al. 2021). As the intrinsic galaxy shape dispersion and the galaxy density vary anisotropically, the shape noise of the sample is no longer determined by the mean value per tomographic bin. Instead, the shape noise becomes an anisotropic distribution too which in the case of KiDS-1000 is skewed below the mean (see Appendix C). In addition, as can be seen by eye from Fig. 7 and 8, variable depth can add angular correlations to the data at the scales of pixels, pointing overlaps, whole pointings or even over the whole footprint, while also changing the signal along the line of sight anisotropically as shown in Fig. 10. These additional correlations can significantly alter the covariance of the two-point statistic or may even lead to non-Gaussian noise.
@@
Figure 16: Comparison of likelihood models with a noise-free mock data vector from the anisotropic systematics forward model. Shown are marginalised posterior contours: in blue, assuming the anisotropic systematics model within the standard setup of KiDS-SBI as described in Sect. 5.4; in black, assuming the same anisotropic systematics forward simulations to train a Mixture Density Network (MDN; Bishop 1994) which is made up of a single multivariate Gaussian; and in pink, applying a standard analysis which assumes a Gaussian likelihood characterised by a numerical sample covariance derived from anisotropic systematics forward simulations and a model given by the analytic pseudo-Cl model described in Appendix D. All posteriors are evaluated assuming the same noise realisation and the same model parameters (see Table 21).
The other distinguishing systematic between the anisotropic systematics model and the standard isotropic systematics model is the shear bias. When including additive and PSF shear biases in the anisotropic systematics model, which can also add a per-cent level (up to ${\sim}0.3\sigma$, but less than $0.1\sigma$ at most angular scales) non-cosmological signal to the pseudo-Cls, mostly driven by the PSF variation shear bias (see Appendix C for details). This is consistent with the negligible impact on the measurement found in previous studies (Giblin et al., 2021). At the same time, the inclusion of the $\alpha_{1}^{(p)}$, $\alpha_{2}^{(p)}$, $c_{1}^{(p)}$ and $c_{2}^{(p)}$ parameters for each tomographic bin and their associated uncertainties adds additional variance to the anisotropic systematics model when we pre-marginalise over these 20 parameters. Therefore, the additive and PSF variation shear biases are yet another source of additional non-cosmological signal and variance.
@@
To assess the effect of model misspecification and whether the inclusion or exclusion of variable depth and shear biases can bias the SBI of cosmological parameters, we repeat the inference of a mock data vector, but this time we generate a random realisation from the standard isotropic systematics model and analyse it using the likelihood learnt from the forward simulation which assumes the anisotropic systematics model, and vice versa. This allows us to determine whether the inclusion of the effects in the forward model would bias our inference if they were not actually present in the data that is analysed. We also show these posteriors in Fig. 15 where we see that when the data model and the likelihood model are mismatched, the constraints are consistent within a given model. This confirms that, given the learned likelihood, the bias in the two-point shear signal from both variable depth and PSF shear bias is negligible in KiDS-1000. These findings are consistent with previous work which found this to be the case for KiDS-1000 for variable depth (Heydenreich et al., 2020; Joachimi et al., 2021) and for the PSF shear bias (Giblin et al., 2021).